Earned Value

1. Complete the calculations

Case

BCWS or PV

ACWP or AC

BCWP or EV

CV

SV

CPI

SPI

1

800

800

800

2

800

600

400

3

800

400

600

4

800

600

600

5

800

800

600

6

800

800

1000
Answer #1

A C D E F G H Case PV AC EV CV SV PI SPI 2 1 =D2/B2 =D3/ B3 800 800 -D2-C2 - D2 / C2 =D2 - B2 800 3 2 800 =D3 / C3 600 400 -D

A C D F G H 1 Case PV AC EV CV SV CPI SPI 2 1 800 800 800 0 0 1 1 3 2 800 600 400 -200 -400 0.67 0.5 4 3 800 400 600 200 -200